OpenStudy (anonymous):

How do you get the inverse of this function 3^x

ganeshie8 (ganeshie8):

\(y = 3^x\) solve x

OpenStudy (anonymous):

x=3^y

OpenStudy (anonymous):

x^3=y

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

is it write?

ganeshie8 (ganeshie8):

you need to take log both sides

ganeshie8 (ganeshie8):

\(\large y = 3^x\) take \(\log_3 \) both sides \(\large \log_3 y = \log_3 3^x\)

ganeshie8 (ganeshie8):

see if u can do the rest

OpenStudy (anonymous):

y= log_3/log_3 3^x f^-1=log_3/log_3 3^x

ganeshie8 (ganeshie8):

\(\large y = 3^x\) take \(\log_3 \) both sides \(\large \log_3 y = \log_3 3^x\) \(\large \log_3 y = x \log_3 3\) \(\large \log_3 y = x \) switch x and y \(\large \log_3 x = y \) \(\large y = \log_3 x \)

ganeshie8 (ganeshie8):

you need to knw few log rules to be able to work above :)

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

np :)